OpenStudy (anonymous):

The volume of a cube is 48 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (mathstudent55):

Here you need the formulas of the volume of a cube and a pyramid.

OpenStudy (anonymous):

Pyramid : V= 1/3 (b)(h) Cube : V= a^3?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Ok, great. Let's draw a cube and a pyramid and compare them. Since a cube has to have a square base, and the pyramid fits inside, the pyramid also has to have a square base.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):